# RESOLVE-IPD: High-Fidelity Individual Patient Data Reconstruction and Uncertainty-Aware Subgroup Meta-Analysis

### v1

@@ Abstract
Purpose
Individual patient data (IPD) from oncology trials are essential for reliable evidence synthesis but are rarely publicly available, necessitating reconstruction from published Kaplan–Meier (KM) curves. Existing reconstruction methods suffer from digitization errors, unrealistic uniform censoring assumptions, and the inability to recover subgroup-level IPD when only aggregate statistics are available. We developed RESOLVE-IPD, a unified computational framework that enables high-fidelity IPD reconstruction and uncertainty-aware subgroup meta-analysis to address these limitations.
Methods
RESOLVE-IPD comprises two components. The first component, High-Fidelity IPD Reconstruction, integrates the VEC-KM and CEN-KM modules: VEC-KM extracts precise KM coordinates and explicit censoring marks from vectorized figures, minimizing digitization error, while CEN-KM corrects overlapping censor symbols and eliminates the uniform censoring assumption. The second component, Uncertainty-Aware Subgroup Recovery, employs the MAPLE (Marginal Assignment of Plausible Labels and Evidence Propagation) algorithm to infer patient-level subgroup labels consistent with published summary statistics (e.g., hazard ratio, median overall survival) when subgroup KM curves are unavailable. MAPLE generates ensembles of mathematically valid labelings, facilitating a propagating meta-analysis that quantifies and reflects uncertainty from subgroup reconstruction.
Results
RESOLVE-IPD was validated through a subgroup meta-analysis of four trials in advanced esophageal squamous cell carcinoma, focusing on the programmed death ligand 1 (PD-L1)-low population. The first-stage modules reconstructed overall survival IPD that precisely matched published KM curves and reported summary statistics. MAPLE successfully inferred subgroup labels and produced plausible IPD reconstructions consistent with reported hazard ratios and survival estimates. Propagating meta-analysis confirmed a survival benefit for immunotherapy over chemotherapy, with significant effects during the 6- to 12-month period.
Conclusion
RESOLVE-IPD enables accurate IPD reconstruction and robust, uncertainty-aware subgroup meta-analyses, strengthening the reliability and transparency of secondary evidence synthesis in precision oncology.
@@ Introduction
The accuracy and utility of such secondary analyses are fundamentally constrained by limitations in existing IPD reconstruction methods [5, 6, 7, 8]. These limitations stem from two main sources: errors in data acquisition and structural flaws in the underlying statistical algorithms. First, traditional approaches rely on manual digitization or low-resolution raster images, introducing coordinate extraction errors. For instance, tools like IPDfromKM [6] require manual calibration, which is both labor-intensive and error-prone. Although newer methods such as KM-GPT [9] employ computer vision for automated coordinate detection, residual inaccuracies persist. Second, the dominant reconstruction algorithm, namely the modified iterative Kaplan–Meier (iKM) method, assumes a uniform censoring mechanism within intervals, as it does not account for censoring marks during reconstruction. This assumption is clinically unrealistic, as censoring usually occurs at scheduled assessments or due to patient dropout, leading to biased survival estimates. While the KMtoIPD algorithm [8] introduced explicit censoring times into reconstruction, it overlooks the issue of overlapping censor marks, in which a single symbol can correspond to multiple censored patients. This limitation results in underestimation of censoring frequency and degraded reconstruction fidelity, particularly under high censoring. These inaccuracies propagate through all subsequent analyses, systematically biasing the estimation of key clinical endpoints such as the hazard ratio and restricted mean survival time (RMST).
@@
To address these methodological and clinical challenges, we develop RESOLVE-IPD: a unified computational framework for high-fidelity IPD reconstruction and robust downstream evidence synthesis, with a particular focus on uncertainty-aware subgroup meta-analyses. RESOLVE-IPD addresses critical limitations in two key areas. First, it resolves the accuracy bottleneck by combining high-fidelity vector graphics extraction with a novel IPD reconstruction algorithm designed to utilize explicit censoring marks, thereby eliminating the systemic bias of the uniform censoring assumption. Second, to recover unreported subgroup information, we introduce an optimization-based method that leverages published summary statistics to infer missing subgroup survival curves from the overall reconstructed IPD. This methodology characterizes the set of mathematically valid, non-unique subgroup label assignments and develops rigorous techniques to quantify the resulting uncertainty, report it, and systematically propagate it into downstream meta-analyses. Collectively, RESOLVE-IPD establishes a new standard for secondary data analysis in oncology by ensuring statistically robust IPD recovery and enabling novel, biomarker-stratified subgroup analyses with fully quantified uncertainty.
@@ Methods
Since this inverse problem is inherently non-identifiable, summary statistics often admit many distinct yet mathematically valid subgroup labelings. To address this, MAPLE first finds configurations with the smallest recovery error, denoted $\mathcal{G}_{\mathrm{MAPLE}}$. When publications include subgroup KM plots only as low-fidelity raster images (e.g., in supplementary materials), the algorithm leverages existing tools like KM-GPT [9] to reconstruct approximate IPD. MAPLE then uses this low-fidelity data to refine the candidate set, filtering labelings to retain only those whose derived subgroup KM curves fall within the 95% CI of the reconstructed curve, denoted $\mathcal{G}_{\mathrm{MAPLE}}^{*}$. This ensures alignment with all available evidence while acknowledging and preserving the inherent uncertainty stemming from non-identifiability.
The output of MAPLE is an ensemble of plausible subgroup labelings, denoted denoted $\mathcal{G}_{\mathrm{MAPLE}}$ or $\mathcal{G}_{\mathrm{MAPLE}}^{*}$ when further refined. Each member of the ensemble represents a distinct yet valid reconstruction of subgroup-level IPD, offering a principled quantification of uncertainty in subgroup assignment. This ensemble provides a principled quantification of uncertainty in subgroup assignment and forms the basis for uncertainty-aware downstream analyses. Detailed optimization formulations and algorithmic implementations of MAPLE are outlined in Supplemental Material Section S2.2.
@@
This approach provides a transparent and quantitative measure of evidence robustness. A narrow distribution indicates that the meta-analytic conclusions are stable despite label uncertainty, whereas a wide distribution reveals sensitivity to subgroup ambiguity, an important consideration for evidence grading and clinical guideline development. By rigorously quantifying and propagating these uncertainties, RESOLVE-IPD strengthens the statistical validity and credibility of meta-analytic evidence in oncology.
@@ Discussion
Code Availability
The implementation code for RESOLVE-IPD, including VEC-KM, CEN-KM, and MAPLE modules, is publicly available at https://github.com/JackZhao0312/RESOLVE-IPD.
@@
This methodology ensures consistency with all available evidence while explicitly preserving the uncertainty inherent in this non-identifiable inverse problem. The final output is an ensemble $\mathcal{G}_{\mathrm{MAPLE}}$ or $\mathcal{G}_{\mathrm{MAPLE}}^{*}$ of plausible subgroup labelings, where each member represents a distinct, mathematically valid reconstruction of the subgroup-level IPD. This ensemble provides principled uncertainty quantification for subgroup assignment and serves as the foundation for uncertainty-aware downstream analyses within the RESOLVE-IPD framework.

### v2

@@ Abstract
Individual patient data (IPD) from oncology trials are essential for reliable evidence synthesis but are rarely publicly available, necessitating reconstruction from published Kaplan-Meier (KM) curves. Existing reconstruction methods suffer from digitization errors, unrealistic uniform censoring assumptions, and the inability to recover subgroup-level IPD when only aggregate statistics are available. To address these limitations, we developed RESOLVE-IPD, a unified computational framework comprising two primary components: high-fidelity IPD reconstruction and uncertainty-aware subgroup recovery. The first component integrates the VEC-KM and CEN-KM modules; VEC-KM extracts precise coordinates and explicit censoring marks from vectorized figures to minimize digitization error, while CEN-KM resolves overlapping censor symbols and eliminates the unrealistic assumption of uniform censoring. The second component utilizes the MAPLE module to identify an ensemble of data-compatible labelings that minimize discrepancies between reconstructed and reported statistics, enabling a meta-analysis framework that explicitly quantifies and propagates the uncertainty inherent in subgroup reconstruction. RESOLVE-IPD was validated through a subgroup meta-analysis of four trials in advanced esophageal squamous cell carcinoma, focusing on the programmed death ligand 1 (PD-L1)-low population. The first-stage modules reconstructed overall survival IPD that precisely matched published KM curves and reported summary statistics. MAPLE successfully inferred subgroup labels and produced plausible IPD reconstructions consistent with reported hazard ratios and survival estimates. Propagating meta-analysis confirmed a survival benefit for immunotherapy over chemotherapy, with significant effects during the 6- to 12-month period.
@@ Introduction
The accuracy and utility of such secondary analyses are fundamentally constrained by limitations in existing IPD reconstruction methods [5, 6, 7, 8, 9]. These limitations stem from two main sources: errors in data acquisition and structural flaws in the underlying statistical algorithms. First, traditional approaches rely on manual digitization or low-resolution raster images, introducing coordinate extraction errors. For instance, tools like IPDfromKM [6] require manual calibration, which is both labor-intensive and error-prone. Although newer methods like KM-GPT [10] leverage computer vision for automated coordinate detection, residual inaccuracies remain. Second, the dominant reconstruction algorithm, namely the modified iterative Kaplan-Meier (iKM) method, assumes a uniform censoring mechanism within intervals, as it does not account for censoring marks during reconstruction [6]. This assumption is clinically unrealistic, as censoring usually occurs at scheduled assessments or due to patient dropout, leading to biased survival estimates. The KMtoIPD algorithm [8] introduced an improvement by allowing digitized censoring times to be incorporated directly into the reconstruction, but it overlooks the issue of overlapping censor marks, in which a single symbol can correspond to multiple censored patients. This limitation results in underestimation of censoring frequency and degraded reconstruction fidelity, particularly under high censoring. More recently, SynthIPD [9] digitizes KM curves at vector-level precision and extracts event and censoring coordinates. Nevertheless, its IPD reconstruction step still estimates the number of events within each risk-table–aligned interval using an allocation approach similar to that of the modified iKM algorithm. As a result, when the risk table is sparse or the censoring rate is high, the interval-specific event estimates may be less reliable, which can limit the accuracy of the reconstructed IPD. These inaccuracies propagate through all subsequent analyses, systematically biasing the estimation of key clinical endpoints such as the hazard ratio and restricted mean survival time (RMST).
@@
To address these methodological and clinical challenges, we develop RESOLVE-IPD: a unified computational framework for high-fidelity IPD reconstruction and robust downstream evidence synthesis, with a particular focus on uncertainty-aware subgroup meta-analyses. RESOLVE-IPD addresses critical limitations in two key areas. First, it resolves the accuracy bottleneck by combining high-fidelity vector graphics extraction with a novel IPD reconstruction algorithm designed to utilize explicit censoring marks, thereby eliminating the systemic bias of the uniform censoring assumption. Second, to recover unreported subgroup information, we introduce an optimization-based method that leverages published summary statistics to infer missing subgroup survival curves from the overall reconstructed IPD. This methodology identifies the set of data-compatible subgroup label assignments, defined as those consistent with the published summary statistics, and develops rigorous techniques to quantify, report, and systematically propagate the resulting uncertainty into downstream meta-analyses. Collectively, RESOLVE-IPD establishes a new standard for secondary data analysis in oncology by ensuring statistically robust IPD recovery and enabling novel, biomarker-stratified subgroup analyses with fully quantified uncertainty.
@@ Methods
Since this inverse problem is inherently non-identifiable, summary statistics often admit many distinct yet data-compatible subgroup labelings. To address this, MAPLE first finds configurations with the smallest recovery error, denoted $\mathcal{G}_{\mathrm{MAPLE}}$. When publications include subgroup KM plots only as low-fidelity raster images (e.g., in supplementary materials), the algorithm leverages existing tools like KM-GPT [10] to reconstruct approximate IPD. MAPLE then uses this low-fidelity data to refine the candidate set, filtering labelings to retain only those whose derived subgroup KM curves fall within the 95% CI of the reconstructed curve, denoted $\mathcal{G}_{\mathrm{MAPLE}}^{*}$. This ensures alignment with all available evidence while acknowledging and preserving the inherent uncertainty stemming from non-identifiability.
The output of MAPLE is an ensemble of plausible subgroup labelings, denoted $\mathcal{G}_{\mathrm{MAPLE}}$ or $\mathcal{G}_{\mathrm{MAPLE}}^{*}$ when further refined. Each member of the ensemble represents a distinct yet data-compatible reconstruction of subgroup-level IPD, offering a principled quantification of uncertainty in subgroup assignment. This ensemble provides a principled quantification of uncertainty in subgroup assignment and forms the basis for uncertainty-aware downstream analyses. Detailed optimization formulations and algorithmic implementations of MAPLE are outlined in Supplemental Material Section S2.2.
@@
This approach provides a transparent and quantitative measure of evidence robustness. A narrow distribution indicates that the meta-analytic conclusions are stable despite label uncertainty, whereas a wide distribution reveals sensitivity to subgroup ambiguity, an important consideration for evidence grading and clinical guideline development. By rigorously quantifying and propagating these uncertainties, RESOLVE-IPD strengthens the statistical validity and credibility of meta-analytic evidence in oncology. The implementation code for RESOLVE-IPD, including VEC-KM, CEN-KM, and MAPLE modules, applied to simulation data, is publicly available at: https://github.com/JackZhao0312/RESOLVE-IPD. The CEN-KM module is also available via a web-based interface at https://km-gpt.wse.jhu.edu/.
@@ Discussion
Acknowledgment
Research reported in this publication was supported by the National Institutes of Health under grants R01MH128085 and R01AI197147.
@@ Acknowledgment
This methodology ensures consistency with all available evidence while explicitly preserving the uncertainty inherent in this non-identifiable inverse problem. The final output is an ensemble $\mathcal{G}_{\mathrm{MAPLE}}$ or $\mathcal{G}_{\mathrm{MAPLE}}^{*}$ of plausible subgroup labelings, where each member represents a distinct, data-compatible reconstruction of the subgroup-level IPD. This ensemble provides principled uncertainty quantification for subgroup assignment and serves as the foundation for uncertainty-aware downstream analyses within the RESOLVE-IPD framework.